Calcium carbide reacts with water to produce acetylene gas. A chemist in a quality control lab is testing the purity of calcium carbide. She measures 9.37 g of grey solid from a tin labelled “calcium carbide”, and reacts it with water to collect 935 mL of acetylene gas at 99.5 kPa and 17.0 °C. What percentage of her original grey solid was calcium carbide?
CaC2(s) + 2 H2O(l) → C2H2(g)
1
Expert's answer
2020-06-22T09:17:56-0400

CaC2+2H2O=C2H2+Ca(OH)2

1.The amount of C2H2 can be found by the following Mendeleev-Clapeyron equation:

n=P*V/R*T where R is equal to 8.31

n=99.5*0.935/8.31*290 =0.0386mol

Comment: 290 is the temperature in Kelvin. To find this, 273 degrees Celsius is added

2.According to the above reaction, 0.0386 mol of C2H2 is formed from as many moles of CaC2. Hence 0.0386mol(2.4704gr) CaC2 was reacted

3.determine the percentage of CaC2 unit in the starting material. To do this, the mass of CaC2 is divided by the mass of the total substance and multiplied by 100%

2.47/9.37=0.26365*100%=26.365%

Answer: 26.365%